# LoRAverse: A Submodular Framework to Retrieve Diverse Adapters for Diffusion Models

### v1

@@ 12 Details of Concept Extractor
A complete example prompt for the concept extractor is provided in the first column of Table 4, utilizing the gpt-4o-mini model. We design a structured prompt to help the LLM effectively identify distinct, orthogonal concepts while merging those that are semantically similar. This module is implemented using LangChain. [5].
@@ 13 Details of Adapter Safety Checker
The decision to filter anthropomorphic content stems from observations that these styles often blur fantasy and reality in problematic ways. These figures frequently appear in exaggerated forms that may contribute to objectification of human-like features. Excluding these adapters by default helps maintain more neutral outputs unless specifically requested. Implemented using LangChain [5], the safety checker introduces computational overhead but remains a practical component. This approach balances creative diversity with safety guards, reducing the risk of generating unsafe outputs while preserving quality and diversity in the results.
@@ 14 Details of VLM-as-a-Judge
The complete prompt used for evaluating the diversity, quality, and textual alignment of the image sets with gpt-4o is outlined in Table 5. Our methodology involves presenting three distinct image sets and using multi-turn prompting techniques to differentiate between them. Each set is numbered sequentially, and the VLM is asked to evaluate their diversity, quality, and textual alignment. The prompt includes a structured rubric, clear instructions, reminders, and example model outputs. To quantify the metrics, the model uses Chain-of-Thought reasoning to assign scores ranging from 0 to 2, based on methodologies from Stylus [9, 40, 42].
@@ 15 Details of Clustering Implementation
To enhance reproducibility, we provide a detailed overview of our clustering methodology. We use the BERTopic [13] framework with UMAP [28] and HDBSCAN [27] for adapter grouping. The UMAP configuration is set to 5 neighbors, 5 components, and cosine distance metric, prioritizing local structure preservation while reducing dimensionality efficiently. For clustering, HDBSCAN is configured with a minimum cluster size of 3, allowing for the formation of micro-clusters. We use euclidean distance in the low-dimensional space and enable prediction data to support unseen data inference. The number of clusters is not predetermined but emerges dynamically from the density-based clustering, ensuring adaptability across datasets. Additionally, we set the number of topics to 10. These design choices capture the underlying structure of LoRA models.
@@ 17 Compatibility with Other Text-to-Image Models
LoRAverse is inherently backbone‑agnostic because it represents each adapter solely by its CLIP embedding, the submodular retrieval objective–and thus the entire pipeline–remains unchanged when the underlying text‑to‑image model is swapped. To validate this claim, we indexed 300 publicly released Flux-compatible LoRA adapters from Hugging Face [1], created their embeddings, and ran our selection algorithm using Flux.1 backbone (see Fig. 11).

### v2

@@ 12 Details of Concept Extractor
A complete example prompt for the concept extractor is provided in the first column of Table 4, utilizing the gpt-4o-mini model. We design a structured prompt to help the LLM effectively identify distinct, orthogonal concepts while merging those that are semantically similar. This module is implemented using LangChain. [Chase_LangChain_2022].
@@ 13 Details of Adapter Safety Checker
The decision to filter anthropomorphic content stems from observations that these styles often blur fantasy and reality in problematic ways. These figures frequently appear in exaggerated forms that may contribute to objectification of human-like features. Excluding these adapters by default helps maintain more neutral outputs unless specifically requested. Implemented using LangChain [Chase_LangChain_2022], the safety checker introduces computational overhead but remains a practical component. This approach balances creative diversity with safety guards, reducing the risk of generating unsafe outputs while preserving quality and diversity in the results.
@@ 14 Details of VLM-as-a-Judge
The complete prompt used for evaluating the diversity, quality, and textual alignment of the image sets with gpt-4o is outlined in Table 5. Our methodology involves presenting three distinct image sets and using multi-turn prompting techniques to differentiate between them. Each set is numbered sequentially, and the VLM is asked to evaluate their diversity, quality, and textual alignment. The prompt includes a structured rubric, clear instructions, reminders, and example model outputs. To quantify the metrics, the model uses Chain-of-Thought reasoning to assign scores ranging from 0 to 2, based on methodologies from Stylus [dunlap2024describingdifferencesimagesets, wei2023chainofthoughtpromptingelicitsreasoning, yao2023treethoughtsdeliberateproblem].
@@ 15 Details of Clustering Implementation
To enhance reproducibility, we provide a detailed overview of our clustering methodology. We use the BERTopic [grootendorst2022bertopic] framework with UMAP [mcinnes2020umapuniformmanifoldapproximation] and HDBSCAN [Malzer_2020] for adapter grouping. The UMAP configuration is set to 5 neighbors, 5 components, and cosine distance metric, prioritizing local structure preservation while reducing dimensionality efficiently. For clustering, HDBSCAN is configured with a minimum cluster size of 3, allowing for the formation of micro-clusters. We use euclidean distance in the low-dimensional space and enable prediction data to support unseen data inference. The number of clusters is not predetermined but emerges dynamically from the density-based clustering, ensuring adaptability across datasets. Additionally, we set the number of topics to 10. These design choices capture the underlying structure of LoRA models.
@@ 17 Compatibility with Other Text-to-Image Models
LoRAverse is inherently backbone‑agnostic because it represents each adapter solely by its CLIP embedding, the submodular retrieval objective–and thus the entire pipeline–remains unchanged when the underlying text‑to‑image model is swapped. To validate this claim, we indexed 300 publicly released Flux-compatible LoRA adapters from Hugging Face [hf_website], created their embeddings, and ran our selection algorithm using Flux.1 backbone (see Fig. 11).